OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values below?

OpenStudy (anonymous):

It is supposed to be a table: x : -2, 0, 4 f(x): 0, -6, 78

OpenStudy (anonymous):

ChoiceS: y=6x^2+5x-4 y=-4x^2-5x+6 y=5x^2+4x-6 y=4x^2+5x-6